OpenStudy (anonymous):

can anyone explain and help me with this? (5x-4)^3 = 64

OpenStudy (anonymous):

Well what do you think you have to do?

OpenStudy (anonymous):

@blahblah_who_cares using square roots to solve?

OpenStudy (anonymous):

Are you sure its square roots?

OpenStudy (anonymous):

no Im not

OpenStudy (anonymous):

can we use roots to solve?

OpenStudy (anonymous):

Well your equation is raised to 3rd power... Hint

OpenStudy (anonymous):

yeah so its not squares then

OpenStudy (anonymous):

how would we solve with roots?

OpenStudy (anonymous):

So what do you think you have to do?

OpenStudy (anonymous):

If its not squaring it, then you would have to......

OpenStudy (anonymous):

Im not sure because its cubed and not squared

OpenStudy (anonymous):

HINT: 64 = 4³

OpenStudy (anonymous):

5x-4=8

OpenStudy (anonymous):

Ok. So if you are not squaring it and ur equation is cubed, you would have to..... hint:you are literally saying it.

OpenStudy (anonymous):

So you are saying 8^3=64?

OpenStudy (anonymous):

cubing it?

OpenStudy (anonymous):

8² = 64 = 4³

OpenStudy (anonymous):

5x-4 \(\neq\) 8

OpenStudy (anonymous):

5x-4=4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4x4x4=64?

OpenStudy (anonymous):

Yes. I told you. 4³ = 64

OpenStudy (anonymous):

4³ = 4·4·4 = 64

OpenStudy (anonymous):

but the part thats confusing is 5x-4=4 and 5x-4=-4, then would how do I write the third part?

OpenStudy (anonymous):

Don't worry about positive or negative, it's only for even exponent. You are dealing with odd exponent.

OpenStudy (anonymous):

ok so then its just 5x-4=4 and 5x-4=-4

OpenStudy (anonymous):

Then solve

OpenStudy (anonymous):

Woah your negatives or positives stay the same because it is to an odd power. ex. -1^3 = -1*-1*-1. two negatives cancel leaving one.

OpenStudy (anonymous):

so the sol set is 0.625, and 0?

OpenStudy (anonymous):

or 5/8 and 0

OpenStudy (anonymous):

No, you are dealing with odd power, thus you only have one answer.

OpenStudy (anonymous):

"your negatives or positives stay the same because it is to an odd power. ex. -1^3 = -1*-1*-1. two negatives cancel leaving one." - blahblah_who_cares

OpenStudy (anonymous):

@micahwood50 so then my answer is5/8?

OpenStudy (anonymous):

Incorrect, but you are close.

OpenStudy (anonymous):

@micahwood50 alright let me try again

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

@micahwood50 can you help me out?

OpenStudy (anonymous):

\[(5x-4)^3 = 64\]\[\sqrt[3]{(5x-4)^3} = \sqrt[3]{64}\]\[5x - 4 = 4\] See what I did there? Can you solve for the x now?

OpenStudy (anonymous):

5x/5=8/4

OpenStudy (anonymous):

x=4?

OpenStudy (anonymous):

8/5*

OpenStudy (anonymous):

sorry 8/5

OpenStudy (anonymous):

^correct

OpenStudy (anonymous):

You divide both sides by 5.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah, it is 8/5

OpenStudy (anonymous):

Is it clear?

OpenStudy (anonymous):

I must be really tired becauase thats the setup I did up there earlier. But just divided incorrectly aggggghhhhh

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

It's alright. I make silly mistakes sometime. Don't worry about it. Glad we helped.

OpenStudy (anonymous):

Everyone makes mistakes.

OpenStudy (anonymous):

“No one is perfect... that's why pencils have erasers.” :P

OpenStudy (anonymous):

Lol thank you!